HYDROGEN INTERACTION WITH METAL HALIDES: THE NUCLEAR QUADRUPOLE COUPLING CONSTANT OF GOLD IN THE p-H_{2}-AuCl COMPLEX AND TRENDS IN THE OTHER HYDROGEN-COINAGE METAL HALIDE INTERACTIONS

\underline{DANIEL A. OBENCHAIN}, G. S. GRUBBS II, HERBERT M. PICKETT, and STEWART E. NOVICK, Department of Chemistry, Wesleyan Univeristy, 52 Lawn Avenue, Middletown, CT, 06459-0180, USA.

The rotational spectrum of p-H_{2}-AuCl has been measured using a laser ablation equipped FTMW cavity spectrometer. The predicted structure, using a 60 electron core potential and an aug-cc-pVQZ^{a}^{b} at the MP2 level of theory, shows that H_{2} has an r_{e} = 0.91Å. A predicted value for the eQq(χ_{aa}) of gold required a semi-empirical method using the results of previous AuCl complexes in the gas phase. Transitions have been measured across multiple J levels, and have been used to determine the rotational constants, centrifugal distortion constants, and nuclear quadrupole coupling constants of multiple isotopologues. The o-H_{2}-AuCl has also been observed. While the monomer value of the eQq of ^{197}Au was determined to be 9.63112(13) MHz in ^{197}Au^{35}Cl^{c}, we observed a significant change in the eQq of ^{197}Au in p-H_{2}-^{197}Au^{35}Cl to a value of -817.9983(36) MHz [0.02728556(12) cm^{−1}], giving rise to a Au hyperfine component splittings of approximately 220 MHz [0.007 cm^{−1}] and suggesting a change in electronic structure with the interaction of H_{2}. We will present a summary of the hydrogen-coinage metal halide talks given at this conference, including trends in eQq and hydrogen metal halide dissociation energies.

aD. Figgen, G. Rauhut, M. Dolh, H. Stoll J. Chem. Phys. 311(2005) 227.

bK. A. Peterson, C. Puzzarini, Theor. Chem. Acc. 114 (2005) 283.

cC. J. Evans, C. L. Gerry J. Mol. Spectrosc. 203 (2000) 105.